OpenStudy (anonymous):

Solve -2log8(x + 1) = -8 pleasee!

OpenStudy (help!!!!):

\[\frac{2\ln \left(1250\right)}{\ln \left(10\right)}\]

OpenStudy (bradely):

-2log8(x + 1) = -8 log8(x + 1)=-8/-2=4 log8(x + 1) =4 x+1=8^4 x+1=4096 x=4095 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

thank you!